• A statistics practitioner wants to test the following hypotheses. H0: mu = 0.70 H1: mu greater than 0.70 A random sample of 100 produced p = 0.73. Calculate the p-value of the test.
1
Expert's answer
2021-07-19T05:52:13-0400

Test statistic: "z=\frac{0.73-0.70}{\sqrt{\frac{0.70(1-0.70)}{100}}}=0.65."

P-value: "p=P(Z>0.65)=0.2578."